# MCHex: Marching Cubes Based Adaptive Hexahedral Mesh Generation with Guaranteed Positive Jacobian

### v1

@@ 1 Introduction
Hex mesh generation has long been regarded as a fundamental yet challenging problem in computational geometry and scientific computing. Hex meshes are highly desirable in finite element analysis (FEA) and isogeometric analysis Zhang (2016) due to their superior numerical properties compared to tetrahedral meshes, offering better convergence rates, lower computational cost, and higher accuracy for many physical simulations Benzley et al. (1995); Cifuentes and Kalbag (1992); Wang et al. (2004, 2021). Despite decades of research, automatically generating high-quality, all-hex meshes that conform to complex geometries while maintaining good mesh quality remains an open problem Schneider et al. (2022); Blacker (2000); Owen (2000); Shepherd and Johnson (2008); Tautges (2001); Zhang (2013). A wide spectrum of approaches has been explored, ranging from advancing front Canann (1992), dual approach Zhang and Bajaj (2006), domain decomposition Armstrong et al. (2015), grid-based Schneiders (2000a), polycube maps Gregson et al. (2011); Yu et al. (2022) to frame fields Nieser et al. (2011); however, most of these methods either require significant user intervention, lack robustness on general inputs, no quality guarantees, or involve long computation time Pietroni et al. (2022).
@@
Among various mesh generation strategies, the grid-based method has become the most reliable approach and the only one widely adopted in industrial practice IntrinSIM (2025); Coreform (2025). The algorithm typically starts by adaptively refining a Cartesian grid until a specified criterion is met. Commonly used criteria include error-sensitive functions Zhang and Bajaj (2006); Zhang et al. (2010, 2013); Hu et al. (2013), normal similarity Ito et al. (2009), local thickness Maréchal (2009); Livesu et al. (2021); Pitzalis et al. (2021); Tong et al. (2024), and surface approximation accuracy Gao et al. (2019); Owen et al. (2017).
@@
After refinement, hanging nodes in the grids are eliminated by replacing cells with transition templates that locally restore mesh conformity. Existing methods can be classified into primal and dual approaches. Primal methods directly incorporate hanging nodes into the output hex mesh. For 3-refinement schemes, the first method was introduced by Schneiders (2000b), but it could not handle concave regions effectively, leading to excessive refinement. Subsequent works improved this by developing templates for concave edges Ito et al. (2009); Elsheikh and Elsheikh (2014). In contrast, 2-refinement templates, though more challenging to implement, can mesh both convex and concave regions with fewer elements Ebeida et al. (2011); Zhang et al. (2013); Owen et al. (2017).
@@
Dual methods, which are exclusively applied to 2-refinement schemes, modify the input grids so that the dual mesh consists entirely of hex elements. These methods represent a significant direction for reducing mesh size. Their primary advantage over primal methods is that transition cells do not propagate across three layers. The development of these methods has been a trajectory of relaxing constraints. The first proposed method required strong balancing and octree pairing Maréchal (2009), a requirement retained in two later pipelines Gao et al. (2019); Tong et al. (2024). Subsequent studies optimized the transition templates to reduce element counts Hu et al. (2013). A significant advancement was made by Livesu et al. (2021), who introduced rotation-symmetric templates to reduce irregular valence and expanded the template set to relax the strong balancing condition to weak balancing and greatly reduced the element count. The most element-efficient method to date further relaxed octree pairing to generalized pairing by using linear programming to ensure an even number of same-density cells in all three directions Pitzalis et al. (2021). It is noteworthy, however, that the reduction in element count does not come without cost. It is achieved at the expense of lowering the upper bound on element quality. For example, based on the authors’ experiments, the original 3-refinement templates Schneiders (2000b) yield a minimum scaled Jacobian of 0.48 without any optimization, while more recent methods employing weak balancing and generalized pairing produce minimum scaled Jacobians on the order of $10^{-3}$ Livesu et al. (2021); Pitzalis et al. (2021). Although optimization techniques can mitigate these issues, the increasingly complex topological connections inherently constrain the upper bound of mesh quality that can be achieved through warping points.
@@
The final step in the pipeline involves boundary projection. Due to the axis-aligned nature of grid-based meshing, boundary vertices must be projected onto the target geometry to accurately approximate the input surface. Early methods projected vertices directly without preprocessing, often resulting in poor element quality at the boundary. This degradation stems from a well-documented issue: when multiple faces of a single hex element lie on the boundary, projecting their vertices independently onto the target surface can induce significant distortion. The combination of padding or pillowing techniques addressed this issue by ensuring that each boundary element has only one face to be projected Mitchell and Tautges (1995). This has since become a standard step in nearly all grid-based meshing pipelines. However, the topological variability of surface vertex stars, ideally homeomorphic to a disk but sometimes more complex, requires careful handling, often involving single or multiple pillowed points depending on the configuration Ito et al. (2009); Gao et al. (2019); Tong et al. (2024). Following pillowing, all methods apply optimization or smoothing techniques to project the pillowed vertices onto the target surface while optimizing mesh quality. Notably, while all preceding steps in the pipeline have deterministic time complexity with guaranteed mesh quality bounds, the final projection step remains largely heuristic. In quadrilateral (quad) and tetrahedral mesh generation, methods already exist that guarantee both perfect boundary conformity and superior angle range, while preserving sharp features Liang et al. (2010); Liang and Zhang (2011, 2014). Although some optimization methods in hex meshing ensure a positive Jacobian Lin et al. (2015), such guarantees are of limited practical value, as they do not provide a lower bound on the improvement of geometric fidelity between the initial axis-aligned boundary and the final optimized boundary. The main observation by the authors is that this fundamental limitation is one of the key reasons for the slow progress in hex mesh generation over the past decade. In contrast, tetrahedral mesh generation, backed by rigorous theoretical guarantees from constrained Delaunay tetrahedralization, both in terms of element quality and boundary preservation, has seen wider adoption in commercial applications.
@@
The MCHex method has been tested on all the models in Thingi10K. This includes a large-scale comparison with the prior RO method, analyzing time complexity, minimum scaled Jacobian, Intersection over Union (IoU), and Hausdorff distance ratio (HR) normalized by the bounding box diagonal. Furthermore, a comparative analysis is presented on models where state-of-the-art methods IntrinSIM (2025); Gao et al. (2019) fail to maintain both positive scaled Jacobian and accurate boundary approximation.
@@ 2 Rationale
The foundational concept for the MCHex approach is inspired by work on quad meshing without cleanup operations Rushdi et al. (2017), as summarized in Figure 2. The left three panels illustrate the RO method commonly employed in previous grid-based hex or quad meshing pipelines. This method exhibits two significant limitations: (1) the resulting boundary is axis-aligned, creating substantial deviation from the input geometry, and (2) as a consequence of this poor initial approximation, subsequent optimization induces severe distortion in boundary elements, making them highly susceptible to inversion. To mitigate such inversions, additional pillowing of boundary elements is typically required. In contrast, the right two panels demonstrate the Marching Squares (MS)-based approach. It introduces the concept of midpoint subdivision for boundary-cut squares to generate all-quad elements. Notably, even before projection, the mesh boundary derived from the red intersection points closely approximates the input geometry. Furthermore, this construction guarantees positive Jacobian values for all elements. The close initial fit ensures that subsequent projection introduces minimal distortion to boundary elements, significantly reducing the occurrence of inversion and the need for corrective optimization. The issue of poorly shaped elements was addressed by implementing a vertex repelling strategy: when intersection points approach too closely to grid points, the grid points are repelled to maintain reasonable edge ratios. Thanks to the direct geometric interpretation of quad Jacobian as the sine value of the interior angle, this procedure theoretically guarantees maintained positive Jacobian quality. In addition, extra templates were introduced to preserve sharp geometric features.
@@
However, extending this approach to three dimensions introduces substantial challenges. The same authors proposed a three-dimensional hex meshing variant Awad et al. (2016), but the demonstrated results were three simple geometries: a sphere, a doll, and a cylinder, without providing any guarantees on mesh quality. Furthermore, the method was restricted to uniform grids, unlike the adaptive grid framework employed in the two-dimensional case. Here, three primary difficulties emerge. First, the original formulation failed to recognize that the method’s essence lies in combining MC with midpoint subdivision. Instead, the authors simplified the problem to five basic cases of planar-cube intersections. In practice, the topological configurations of input geometry within a cube are considerably more complex, and intersection points on cutting surfaces with more than three vertices are frequently non-coplanar. Second, while midpoint subdivision of an arbitrary $n$-gon in two dimensions invariably produces $n$ quads, the three-dimensional case imposes stricter constraints. As discussed later in Section 4, only polyhedra with 3-regular graph connectivity containing $n$ vertices are bounded to yield $n$ hexes after midpoint subdivision. Third, the geometric interpretation of Jacobian quality becomes significantly more complex. In two-dimensions, the Jacobian is related to the sine of the interior angles; in three-dimensions, it corresponds to the signed volume of the tetrahedron formed by a vertex’s three edges. This geometric complexity, combined with the diverse configurations in adaptive grids, makes establishing rigorous quality guarantees significantly more difficult. This paper aims to provide a comprehensive and rigorous treatment of these three challenges, presenting theoretical foundations and practical solutions for robust three-dimensional hex meshing.
@@ 3 Algorithm Overview
Of particular interest in the storage of three-dimensional geometry is a closed pure two-dimensional simplicial complex $T_{\text{geom}}$. According to the Jordan-Brouwer separation theorem Hatcher (2001), $T_{\text{geom}}$ partitions $\mathbb{R}^{3}$ into two connected components: a bounded internal region $\text{Int}(T_{\text{geom}})$ and an unbounded external region $\mathbb{R}^{3}\setminus\text{Int}(T_{\text{geom}})$. To address computer memory constraints, an axis-aligned bounding box (AABB) $T_{\text{AABB}}$ is introduced, which also partitions $\mathbb{R}^{3}$ into a bounded internal region $\text{Int}(T_{\text{AABB}})$ and an unbounded external region $\mathbb{R}^{3}\setminus\text{Int}(T_{\text{AABB}})$, with the condition that $\text{Int}(T_{\text{geom}})\subset\text{Int}(T_{\text{AABB}})$.
Let $M\subset\mathbb{R}^{3}$ be a solid region, the hex mesh $H_{M}=(V,E,Q,H)$ is a CW-complex Hatcher (2001) of hex elements $H$ that decomposes $M$. $V,E,Q$ and $H$ represent the vertices, the edges, the quad faces, and the hex elements, respectively. The boundary $\partial H$ consists of those faces in $Q$ belonging to only one hex element, along with their constituent edges and vertices. Hex elements in $H$ that are not part of $\partial H$ are referred to as internal elements. The domain $\text{Int}(T_{\text{AABB}})$ is tessellated into hexes $H_{\text{Int}(T_{\text{AABB}})}$. Given the hex mesh $H_{\text{Int}(T_{\text{AABB}})}$, the algorithm extracts $H_{\text{Int}(T_{\text{geom}})}$ and $H_{\text{Int}(T_{\text{AABB}})\setminus\text{Int}(T_{\text{geom}})}$, both of which are bounded to satisfy the following properties: positive Jacobian, accurate boundary (i.e., correct topology, high IoU, and low HR), and $C^{0}$-conformity along their shared boundary. Specifically, “adaptivity” requires the mesh to be denser near high-frequency features and coarser in smoother regions to save elements. $C^{0}$-conformity” means two piecewise linear meshes are topologically and geometrically consistent across their common quad boundary faces. In detail, there exists a bijection $\phi:\partial H_{\text{Int}(T_{\text{geom}})}\to\partial H_{\text{Int}(T_{\text{AABB}})\setminus\text{Int}(T_{\text{geom}})}$ such that for each quad face $q(V_{q},E_{q})\in\partial H_{\text{Int}(T_{\text{geom}})}$, face $\phi(q)\in\partial H_{\text{Int}(T_{\text{AABB}})\setminus\text{Int}(T_{\text{geom}})}$ satisfies $\phi(v)=v$ and $\phi(e)=e$ for $v\in V_{q}$ and $e\in E_{q}$.
@@
The complete pipeline is summarized in Algorithm 1. The process begins by initializing grids using a 3-refinement strategy. Each cell is recursively subdivided into twenty-seven cells until the feature-sensitive error function for every cell falls below a user-defined threshold. The original feature-sensitive error function, defined as $\sum_{i=1}^{27}\frac{\lvert f^{j+1}(\mathbf{x}_{i})-f^{j}(\mathbf{x}_{i})\rvert}{\lvert\nabla f^{j}(\mathbf{x}_{i})\rvert}$ Zhang et al. (2005), where $\mathbf{x}_{i}$ are position vectors at corners of eight cells after 2-refinement, $j$ is the grid level, has been modified to $\max_{i=1}^{64}\frac{\lvert f^{j+1}(\mathbf{x}_{i})-f^{j}(\mathbf{x}_{i})\rvert}{\lvert\nabla f^{j}(\mathbf{x}_{i})\rvert}$, where $\mathbf{x}_{i}$ are position vectors at corners of twenty-seven cells after 3-refinement, $j$ is the grid level, in order to fit to the 3-refinement scheme Schneiders (2000b) and prioritize the subdivision of cells with extreme values. Subsequently, the standard 3-refinement templates are applied to eliminate hanging nodes, resulting in an adaptive hex mesh, $H_{\text{Int}(T_{\text{AABB}})}$, where the SDF value is computed for every grid point. Next, the MC method is employed within each cell. Since the SDF values at all eight grids are known, an isosurface can be constructed for each cell, dividing the cell into an inside and an outside 3-regular graph. The final step involves subdividing every 3-regular graph using the midpoint subdivision. This subdivision produces an all-hex mesh that partitions the domain into $H_{\text{Int}(T_{\text{geom}})}$ (inside the geometry) and $H_{\text{Int}(T_{\text{AABB}})\setminus\text{Int}(T_{\text{geom}})}$ (outside the geometry), separated by the reconstructed isosurface. To guarantee that all hexes have a positive Jacobian, some volume centroids are warped based on theoretically-derived positions which will be explained later in Section 5.

### v2

@@ 1 Introduction
Hex mesh generation has long been regarded as a fundamental yet challenging problem in computational geometry and scientific computing. Hex meshes are highly desirable in finite element analysis (FEA) and isogeometric analysis (zhang2016geometric) due to their superior numerical properties compared to tetrahedral meshes, offering better convergence rates, lower computational cost, and higher accuracy for many physical simulations (benzley1995comparison, cifuentes1992performance, wang2004back, wang2021comparison). Despite decades of research, automatically generating high-quality, all-hex meshes that conform to complex geometries while maintaining good mesh quality remains an open problem (schneider2022large, blacker2000meeting, owen1998survey, shepherd2008hexahedral, tautges2001generation, zhang2013challenges). A wide spectrum of approaches has been explored, ranging from advancing front (canann1992plastering), dual approach (zhang2006adaptive), domain decomposition (armstrong2015common), grid-based (schneiders2000algorithms), polycube maps (gregson2011all, yu2022hexgen) to frame fields (nieser2011cubecover); however, most of these methods either require significant user intervention, lack robustness on general inputs, no quality guarantees, or involve long computation time (pietroni2022hex).
@@
Among various mesh generation strategies, the grid-based method has become the most reliable approach and the only one widely adopted in industrial practice (MeshGems, CoreformCubit). The algorithm typically starts by adaptively refining a Cartesian grid until a specified criterion is met. Commonly used criteria include error-sensitive functions (zhang2006adaptive, zhang2010automatic, zhang2013robust, hu2013adaptive), normal similarity (ito2009octree), local thickness (marechal2009advances, livesu2021optimal, pitzalis2021generalized, tong2024hybridoctree_hex), and surface approximation accuracy (gao2019feature, owen2017template).
@@
After refinement, hanging nodes in the grids are eliminated by replacing cells with transition templates that locally restore mesh conformity. Existing methods can be classified into primal and dual approaches. Primal methods directly incorporate hanging nodes into the output hex mesh. For 3-refinement schemes, the first method was introduced by (schneiders2000octree), but it could not handle concave regions effectively, leading to excessive refinement. Subsequent works improved this by developing templates for concave edges (ito2009octree, elsheikh2014consistent). In contrast, 2-refinement templates, though more challenging to implement, can mesh both convex and concave regions with fewer elements (ebeida2011isotropic, zhang2013robust, owen2017template).
@@
Dual methods, which are exclusively applied to 2-refinement schemes, modify the input grids so that the dual mesh consists entirely of hex elements. These methods represent a significant direction for reducing mesh size. Their primary advantage over primal methods is that transition cells do not propagate across three layers. The development of these methods has been a trajectory of relaxing constraints. The first proposed method required strong balancing and octree pairing (marechal2009advances), a requirement retained in two later pipelines (gao2019feature, tong2024hybridoctree_hex). Subsequent studies optimized the transition templates to reduce element counts (hu2013adaptive). A significant advancement was made by (livesu2021optimal), who introduced rotation-symmetric templates to reduce irregular valence and expanded the template set to relax the strong balancing condition to weak balancing and greatly reduced the element count. The most element-efficient method to date further relaxed octree pairing to generalized pairing by using linear programming to ensure an even number of same-density cells in all three directions (pitzalis2021generalized). It is noteworthy, however, that the reduction in element count does not come without cost. It is achieved at the expense of lowering the upper bound on element quality. For example, based on the authors’ experiments, the original 3-refinement templates (schneiders2000octree) yield a minimum scaled Jacobian of 0.48 without any optimization, while more recent methods employing weak balancing and generalized pairing produce minimum scaled Jacobians on the order of $10^{-3}$ (livesu2021optimal, pitzalis2021generalized). Although optimization techniques can mitigate these issues, the increasingly complex topological connections inherently constrain the upper bound of mesh quality that can be achieved through warping points.
@@
The final step in the pipeline involves boundary projection. Due to the axis-aligned nature of grid-based meshing, boundary vertices must be projected onto the target geometry to accurately approximate the input surface. Early methods projected vertices directly without preprocessing, often resulting in poor element quality at the boundary. This degradation stems from a well-documented issue: when multiple faces of a single hex element lie on the boundary, projecting their vertices independently onto the target surface can induce significant distortion. The combination of padding or pillowing techniques addressed this issue by ensuring that each boundary element has only one face to be projected (mitchell1995pillowing). This has since become a standard step in nearly all grid-based meshing pipelines. However, the topological variability of surface vertex stars, ideally homeomorphic to a disk but sometimes more complex, requires careful handling, often involving single or multiple pillowed points depending on the configuration (ito2009octree, gao2019feature, tong2024hybridoctree_hex). Following pillowing, all methods apply optimization or smoothing techniques to project the pillowed vertices onto the target surface while optimizing mesh quality. Notably, while all preceding steps in the pipeline have deterministic time complexity with guaranteed mesh quality bounds, the final projection step remains largely heuristic. In quadrilateral (quad) and tetrahedral mesh generation, methods already exist that guarantee both perfect boundary conformity and superior angle range, while preserving sharp features (liang2010guaranteed, liang2011hexagon, liang2014octree). Although some optimization methods in hex meshing ensure a positive Jacobian (lin2015quality), such guarantees are of limited practical value, as they do not provide a lower bound on the improvement of geometric fidelity between the initial axis-aligned boundary and the final optimized boundary. The main observation by the authors is that this fundamental limitation is one of the key reasons for the slow progress in hex mesh generation over the past decade. In contrast, tetrahedral mesh generation, backed by rigorous theoretical guarantees from constrained Delaunay tetrahedralization, both in terms of element quality and boundary preservation, has seen wider adoption in commercial applications.
@@
The MCHex method has been tested on all the models in Thingi10K. This includes a large-scale comparison with the prior RO method, analyzing time complexity, minimum scaled Jacobian, Intersection over Union (IoU), and Hausdorff distance ratio (HR) normalized by the bounding box diagonal. Furthermore, a comparative analysis is presented on models where state-of-the-art methods (MeshGems, gao2019feature) fail to maintain both positive scaled Jacobian and accurate boundary approximation.
@@ 2 Rationale
The foundational concept for the MCHex approach is inspired by work on quad meshing without cleanup operations (rushdi2017all), as summarized in Figure 2. The left three panels illustrate the RO method commonly employed in previous grid-based hex or quad meshing pipelines. This method exhibits two significant limitations: (1) the resulting boundary is axis-aligned, creating substantial deviation from the input geometry, and (2) as a consequence of this poor initial approximation, subsequent optimization induces severe distortion in boundary elements, making them highly susceptible to inversion. To mitigate such inversions, additional pillowing of boundary elements is typically required. In contrast, the right two panels demonstrate the Marching Squares (MS)-based approach. It introduces the concept of midpoint subdivision for boundary-cut squares to generate all-quad elements. Notably, even before projection, the mesh boundary derived from the red intersection points closely approximates the input geometry. Furthermore, this construction guarantees positive Jacobian values for all elements. The close initial fit ensures that subsequent projection introduces minimal distortion to boundary elements, significantly reducing the occurrence of inversion and the need for corrective optimization. The issue of poorly shaped elements was addressed by implementing a vertex repelling strategy: when intersection points approach too closely to grid points, the grid points are repelled to maintain reasonable edge ratios. Thanks to the direct geometric interpretation of quad Jacobian as the sine value of the interior angle, this procedure theoretically guarantees maintained positive Jacobian quality. In addition, extra templates were introduced to preserve sharp geometric features.
@@
However, extending this approach to three dimensions introduces substantial challenges. The same authors proposed a three-dimensional hex meshing variant (awad2016all), but the demonstrated results were three simple geometries: a sphere, a doll, and a cylinder, without providing any guarantees on mesh quality. Furthermore, the method was restricted to uniform grids, unlike the adaptive grid framework employed in the two-dimensional case. Here, three primary difficulties emerge. First, the original formulation failed to recognize that the method’s essence lies in combining MC with midpoint subdivision. Instead, the authors simplified the problem to five basic cases of planar-cube intersections. In practice, the topological configurations of input geometry within a cube are considerably more complex, and intersection points on cutting surfaces with more than three vertices are frequently non-coplanar. Second, while midpoint subdivision of an arbitrary $n$-gon in two dimensions invariably produces $n$ quads, the three-dimensional case imposes stricter constraints. As discussed later in Section 4, only polyhedra with 3-regular graph connectivity containing $n$ vertices are bounded to yield $n$ hexes after midpoint subdivision. Third, the geometric interpretation of Jacobian quality becomes significantly more complex. In two-dimensions, the Jacobian is related to the sine of the interior angles; in three-dimensions, it corresponds to the signed volume of the tetrahedron formed by a vertex’s three edges. This geometric complexity, combined with the diverse configurations in adaptive grids, makes establishing rigorous quality guarantees significantly more difficult. This paper aims to provide a comprehensive and rigorous treatment of these three challenges, presenting theoretical foundations and practical solutions for robust three-dimensional hex meshing.
@@ 3 Algorithm Overview
Of particular interest in the storage of three-dimensional geometry is a closed pure two-dimensional simplicial complex $T_{\text{geom}}$. According to the Jordan-Brouwer separation theorem (hatcher2001algebraic), $T_{\text{geom}}$ partitions $\mathbb{R}^{3}$ into two connected components: a bounded internal region $\text{Int}(T_{\text{geom}})$ and an unbounded external region $\mathbb{R}^{3}\setminus\text{Int}(T_{\text{geom}})$. To address computer memory constraints, an axis-aligned bounding box (AABB) $T_{\text{AABB}}$ is introduced, which also partitions $\mathbb{R}^{3}$ into a bounded internal region $\text{Int}(T_{\text{AABB}})$ and an unbounded external region $\mathbb{R}^{3}\setminus\text{Int}(T_{\text{AABB}})$, with the condition that $\text{Int}(T_{\text{geom}})\subset\text{Int}(T_{\text{AABB}})$.
Let $M\subset\mathbb{R}^{3}$ be a solid region, the hex mesh $H_{M}=(V,E,Q,H)$ is a CW-complex (hatcher2001algebraic) of hex elements $H$ that decomposes $M$. $V,E,Q$ and $H$ represent the vertices, the edges, the quad faces, and the hex elements, respectively. The boundary $\partial H$ consists of those faces in $Q$ belonging to only one hex element, along with their constituent edges and vertices. Hex elements in $H$ that are not part of $\partial H$ are referred to as internal elements. The domain $\text{Int}(T_{\text{AABB}})$ is tessellated into hexes $H_{\text{Int}(T_{\text{AABB}})}$. Given the hex mesh $H_{\text{Int}(T_{\text{AABB}})}$, the algorithm extracts $H_{\text{Int}(T_{\text{geom}})}$ and $H_{\text{Int}(T_{\text{AABB}})\setminus\text{Int}(T_{\text{geom}})}$, both of which are bounded to satisfy the following properties: positive Jacobian, accurate boundary (i.e., correct topology, high IoU, and low HR), and $C^{0}$-conformity along their shared boundary. Specifically, “adaptivity” requires the mesh to be denser near high-frequency features and coarser in smoother regions to save elements. $C^{0}$-conformity” means two piecewise linear meshes are topologically and geometrically consistent across their common quad boundary faces. In detail, there exists a bijection $\phi:\partial H_{\text{Int}(T_{\text{geom}})}\to\partial H_{\text{Int}(T_{\text{AABB}})\setminus\text{Int}(T_{\text{geom}})}$ such that for each quad face $q(V_{q},E_{q})\in\partial H_{\text{Int}(T_{\text{geom}})}$, face $\phi(q)\in\partial H_{\text{Int}(T_{\text{AABB}})\setminus\text{Int}(T_{\text{geom}})}$ satisfies $\phi(v)=v$ and $\phi(e)=e$ for $v\in V_{q}$ and $e\in E_{q}$.
@@
The complete pipeline is summarized in Algorithm 1. The process begins by initializing grids using a 3-refinement strategy. Each cell is recursively subdivided into twenty-seven cells until the feature-sensitive error function for every cell falls below a user-defined threshold. The original feature-sensitive error function, defined as $\sum_{i=1}^{27}\frac{\lvert f^{j+1}(\mathbf{x}_{i})-f^{j}(\mathbf{x}_{i})\rvert}{\lvert\nabla f^{j}(\mathbf{x}_{i})\rvert}$ (zhang20053d), where $\mathbf{x}_{i}$ are position vectors at corners of eight cells after 2-refinement, $j$ is the grid level, has been modified to $\max_{i=1}^{64}\frac{\lvert f^{j+1}(\mathbf{x}_{i})-f^{j}(\mathbf{x}_{i})\rvert}{\lvert\nabla f^{j}(\mathbf{x}_{i})\rvert}$, where $\mathbf{x}_{i}$ are position vectors at corners of twenty-seven cells after 3-refinement, $j$ is the grid level, in order to fit to the 3-refinement scheme (schneiders2000octree) and prioritize the subdivision of cells with extreme values. Subsequently, the standard 3-refinement templates are applied to eliminate hanging nodes, resulting in an adaptive hex mesh, $H_{\text{Int}(T_{\text{AABB}})}$, where the SDF value is computed for every grid point. Next, the MC method is employed within each cell. Since the SDF values at all eight grids are known, an isosurface can be constructed for each cell, dividing the cell into an inside and an outside 3-regular graph. The final step involves subdividing every 3-regular graph using the midpoint subdivision. This subdivision produces an all-hex mesh that partitions the domain into $H_{\text{Int}(T_{\text{geom}})}$ (inside the geometry) and $H_{\text{Int}(T_{\text{AABB}})\setminus\text{Int}(T_{\text{geom}})}$ (outside the geometry), separated by the reconstructed isosurface. To guarantee that all hexes have a positive Jacobian, some volume centroids are warped based on theoretically-derived positions which will be explained later in Section 5.